OpenStudy (anonymous):

When you recieive a loan to make a purchase, you often must make a down payment in cash. The amount of the loan is the purchase cost minus the down payment. Riva made a down payment of $1500 on a used car. She received a loan of $2600. Write and solve an equation to find the cost of the car. Show that your answer is resonable. IM SO LOST!!!! Help i will give badge.

OpenStudy (cwrw238):

cost = down payment + amount of loan c = d + l d = 1500 and l = 2600 dollars c = 1500 + 2600

OpenStudy (anonymous):

so the cost is 4100. right??

OpenStudy (anonymous):

Thanks btw!!! :)

OpenStudy (cwrw238):

yes